How do i find the conic equation for 4(x+1)^2-2(y-3)^2=1
Answer by lwsshak3(11628)

How do i find the conic equation for
4(x+1)^2-2(y-3)^2=1
(x+1)^2/1/4-(y-3)^2/1/2=1
This is an equation of a hyperbola with horizontal transverse axis of the standard form:
(x-h)^2/a^2-(y-k)^2/b^2=1, (h,k) being the (x,y) coordinates of the center.
For given equation:
center: (-1,3)
a^2=1/4
b^2=1/2